OpenStudy (anonymous):

if x+y+z=1 , xy+yx+zx = -1 and xyz=-1 then find the value of x^3+y^3+z^3

OpenStudy (anonymous):

any1 help me

OpenStudy (anonymous):

\[x+y+z = 1\] \[xy+yx+zx = -1\] \[xyz = -1\] this is right?

OpenStudy (anonymous):

yea tanvi that is right

OpenStudy (anonymous):

in the second one, xy+yx = 2xy

OpenStudy (anonymous):

you want to get (x+y+z)^3 and write it out:\[(x+y+z)^3=x^3+y^3+z^3+3(x^2y+x^2z+y^2z+xy^2+xz^2+yz^2)+6xyz\]

OpenStudy (anonymous):

find the value of \[x^3+y^3+z^3\]

OpenStudy (anonymous):

yes joseph sir what after then ?

OpenStudy (anonymous):

that means one of x, y or z is a one. the others can be any twp numbers that are reciprocals of each other.

OpenStudy (anonymous):

then playing around with the terms we get: \[x^3+y^3+z^3+3(x^2y+x^2z+y^2z+xy^2+xz^2+yz^2+3xyz)-3xyz\] \[x^3+y^3+z^3+3(x+y+z)(xy+yz+xz)-3xyz\]

OpenStudy (anonymous):

oh yes and then easily we can put the values i got it joesh sir thanks very much

OpenStudy (anonymous):

sir r u on facebook ?

OpenStudy (anonymous):

so we now have:\[(x+y+z)^3=x^3+y^3+z^3+3(x+y+z)(xy+yz+xz)-3xyz\] just put in the values and solve for your sum.

OpenStudy (anonymous):

Professor Joe

OpenStudy (anonymous):

for completion sake: \[(1)^3 = x^3+y^3+z^3+3(1)(-1)-3(-1) \iff x^3+y^3+z^3=1+3-3 = 1\]

OpenStudy (anonymous):

joseph sir r u on facebook ?

OpenStudy (anonymous):

Im not sure =/ i had an account a looooong time ago, but when i started working at a high school i stopped using it (job rules)

OpenStudy (anonymous):

oh yes u r right sir ...........

OpenStudy (anonymous):

btw, if you wanted the actual solutions to this system, its the same as solving the cubic equation: \[x^3-x^2-x+1=0\]The roots to this equations are the solutions to this system.

OpenStudy (anonymous):

joseph sir help me here : http://openstudy.com/groups/mathematics/updates/4e5f63e70b8b1f45b49de337

OpenStudy (anonymous):

\[x^2(x-1)-(x-1)=0\iff (x^2-1)(x-1)=0\iff (x+1)(x-1)^2=0\] so x=-1, y=1, z =1 (and any permutation) is a solution to the system

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

joe can you help me out

OpenStudy (anonymous):

i can try, whats the question?